# S2C2A: A Flexible Task Space Planning and Control Strategy for Modular Soft Robot Arms

### v1

@@ V-A Position Control
Figure 6: The target (red) trajectories in the position control task and real trajectories applying $NN_{C2A}$ (blue) and CC controller(green).
@@
Figure 7: The MSRA motion in position control using the configuration controller $NN_{C2A}$. The MSRA first decreases the height (1-2) and draws a circle (2-3-4-5). The white dotted lines represent the MSRA end trajectories, and the red arrows represent the motion directions.
@@
Of note, the MSRA states collected by the optical tracking system are only used for comparing with the target trajectory, and we do not utilize them as feedback in the following experiments, which is proven obviously in Section V-E. Only encoder feedback is fed into the configuration controller. The experimental results show that the configuration controller $NN_{C2A}$ outperforms the CC controller; hence, we employ it in the following experiments.
V-B Position and Orientation Control
@@
In addition to the position trajectory shown in Figure 6, we add target angle trajectories for MSRA end orientation, as depicted in the red lines in Figure 9. In these tasks, we include the losses $L_{p},L_{d}$ and the orientation loss $L_{o}$ for tracking. First, we define three tasks requiring the bending angle to 40, 50, and 60, as shown in Figure10-(A), (B), and (C). In the fourth task, the MSRA is controlled to keep the end upward while following a circle, as illustrated in Figure 10-(D). In these tasks, $u_{p}=1$, $u_{d}=0.5$, and $u_{o}=2$. The errors are shown in Table III. All the experiment videos can be found in the Supplementary Video.
@@
Figure 9: The target (red) and real (blue) orientation trajectories in the position and orientation control task for 50.
TABLE III: Errors and standard derivations in position and orientation control
task position error orientation error (z) orientation error (x)
40 3.9$\pm$1.6 cm 3.1$\pm$3.7 /
50 3.2$\pm$1.3 cm 3.9$\pm$3.9 /
60 3.1$\pm$1.5 cm 2.8$\pm$3.3 /
0 2.7$\pm$1.1 cm 4.3$\pm$2.6 /
50a20 3.4$\pm$2.0 cm 1.7$\pm$1.9 4.5$\pm$4.9
50c20 3.4$\pm$1.6 cm 3.9$\pm$3.2 6.0$\pm$4.9
Based on the task 50, we also control the MSRA to bend 20 anticlockwise (50a20) and clockwise (50c20) on the x-y plane. The errors are shown in Table III. We show the MSRA deformation on the same end position in Figure 10-(E), (F), and (G). The MSRA orientations on the x-y plane in these tasks are shown in Figure 8. It is obvious that our approach can perform position and orientation control simultaneously and achieve satisfying accuracy even in some intricate tasks.
V-C Position Constraint
In these tasks, we aim to keep the end position of the base, middle, and end module invariant during motion, as depicted in Figure 11. We include the losses $L_{p},L_{d},L_{o}$ for this task. In task ’base,’ we control the MSRA end to follow a position trajectory while maintaining the base module position static, as shown in Figure 11-(A). In tasks ’middle’ and ’end,’ we aim to keep the middle and end module end positions static and change the same module end orientations, as shown in Figure 11-(B) and (C). The errors are shown in Table IV. During these motions, the corresponding module end positions remain static, but the orientations keep changing for MSRA motion. All the experiment videos can be found in the Supplementary Video.
TABLE IV: Errors and standard derivations in position constraint
@@ V-C Position Constraint
V-D Obstacle Avoidance
In obstacle avoidance, we include the losses $L_{p},L_{d}$ for position control and leverage the obstacle loss $L_{ob}$. First, we set a red sponge bar as the target and control the MSRA end to reach it, as illustrated in Figure 12-(A). Then, we put one blue sponge bar as the obstacle in the last MSRA end trajectory and include $L_{ob}$ for replanning ($u_{ob}=1$). The MSRA end first moves away from the obstacle to avoid collision (Figure 12-(B2)) and then reaches the target (Figure 12-(B3)). Furthermore, we set one more obstacle in the last MSRA end trajectory as depicted in Figure 12-(C). In this case, the S2C planning approach changes the trajectory and helps the MSRA reach the target.
By adjusting the obstacle threshold $r$ mentioned in Equation 5, the MSRA will try to escape the obstacle when it enters the obstacle areas with different sizes. In this case, the obstacles can have different ’risk levels.’ As shown in Figure 13, two obstacles are in the way of the MSRA end to the target. If both obstacles are at the high-risk level, which means the obstacle threshold is large, the MSRA will vibrate near the obstacle boundaries instead of taking risks and rushing to the target, as shown in Figure 13-(B). However, if one obstacle is at low risk, like the left one in Figure 13-(A) and the right one in Figure 13-(C), S2C will generate available trajectories reaching the target and the MSRA end will move to the target. The trajectories will be away from the high-risk obstacles and near the low-risk ones. All the experiment videos can be found in the Supplementary Video.
@@ V-D Obstacle Avoidance
V-E Online Interaction
If the obstacle and target will move during the task, it is necessary to plan the trajectory online. We include the losses $L_{p},L_{d},L_{ob}$ for position control and obstacle avoidance in this task. During the motion, we cover the optical tracker markers with a red cloth to highlight that we achieve accurate control via only motor encoder feedback. As shown in Figure 14-(A), a blue obstacle is moved to reach the MSRA end by hand, and the MSRA end will leave the obstacle as long as it is inside the obstacle area, whose size depends on the obstacle threshold $r$. In addition to online obstacle avoidance, target following can also be fulfilled online. In Figure 14-(B), the MSRA end will follow the moving target. The MSRA end will stop if the target is out of the working space, but it will start to reach the target again if the target reenters the working space, as shown in Figure 14-(C). All the experiment videos can be found in the Supplementary Video.
The online target following motion is steady because the MSRA always aims to minimize the distance between the target and MSRA, which is motivated by the loss $L_{p}$. Meanwhile, in online obstacle avoidance, the MSRA motion is discontinuous. When the obstacle moves near the MSRA and their distance is smaller than $r$ in Equation 5, the MSRA will start to move away from the obstacle until the distance is larger than $r$. At this time, the MSRA will stop because of the configuration change loss $L_{d}$ preventing unnecessary motion, which leads to such discontinuous motion. MSRA will start to move when the obstacle catches the MSRA end again. The obstacle threshold $r$ constraints the obstacle avoidance $L_{ob}$ into a local area, hence the loss will only take effect when the MSRA is within the ’risk area,’ as the bordered blue circles in Figure 13 and 14. Therefore, we endow the MSRA with situation awareness to some extent, which may be significant in the interaction tasks in future work.
@@ V-E Online Interaction
VI Conclusion and Discussion
This work proposes a task space planning and control strategy S2C2A specifically for MSRAs that is flexible to tasks and controllers. We train a biLSTM network $NN_{C2S}$ mapping from configuration to task space as the MSRA forward model and leverage it in an optimization problem S2C for configuration planning. Then, we train a biLSTM network $NN_{C2A}$ for configuration control C2A and compare it with a physical CC controller. Utilizing a cable-driven MSRA, we demonstrate that our approach can achieve simple tasks like position control. Furthermore, some challenging tasks have been managed, such as position and orientation control, obstacle avoidance, and online planning. All the tasks can be achieved with low errors ($<4\%$ in position and $<5^{\circ}$ in orientation).
This flexible planning and control strategy endows MSRAs with accurate and complex deformation capability. Owing to modularity, MSRAs have the potential to perform sophisticated tasks, but effective control approaches specifically for MSRAs have not been fully explored. This work is a benchmark for MSRA control targeted at complicated deformation, enabling MSRA with flexibility and situation awareness. For instance, MSRAs can be utilized for minimally invasive surgery leveraging the position constraint ability and human interaction leveraging online planning ability.
We plan to design better controllers based on S2C2A in future work. For example, internal sensing errors will accumulate along the module sequence. More accurate configuration feedback may be possible by using IMU[61] and flex sensor[62]. It is challenging to estimate whether one target position is feasible for one MSRA, and the difficulty will increase when including target orientation additionally. A novel controller is required to achieve whole shape control, not only just module end state control. Moreover, we would like to take full advantage of the situation awareness ability equipped by our strategy and perform some more intricate interaction tasks.

### v2

@@ IV-B Position Control
Figure 8: The target (red) trajectories in the position control task and real trajectories applying our strategy (blue) and PCC strategy (green).
@@
Figure 9: The MSRA motion in position control using the configuration controller $NN_{C2A}$. The MSRA first decreases the height (0-1) and draws a circle (1-2-3-4-5). The white dotted lines represent the MSRA end trajectories, and the red arrows represent the motion directions.
@@
Of note, the MSRA states collected by the optical tracking system are only used for comparing with the target trajectory, and we do not utilize them as feedback in the following experiments, which is proven obviously in Section IV-F. Only encoder feedback is fed into the configuration controller. The experimental results show that both the PCC approach and our approach can achieve relatively low errors ($\leq 2.5cm$) in position control, and our strategy outperforms the PCC one [2].
IV-C Position and Orientation Control
@@
In addition to the position trajectory shown in Fig. 8, we add target angle trajectories for MSRA end orientation, as depicted in the red lines in Fig. 10. In these tasks, we include the losses $L_{p},L_{d}$ and the orientation loss $L_{o}$ for tracking.
Figure 10: The target (red) and measured (blue) orientation trajectories in the position and orientation control task for the task 50 anticlockwise 20.
First, we define three tasks requiring the bending angle to 40, 50, and 60, as shown in Fig. 11(A), (B), and (C). In the fourth task, the MSRA is controlled to keep the end upward while following a circle, as illustrated in Fig. 11(D). In these tasks, $u_{p}=1$, $u_{d}=0.5$, and $u_{o}=2$. The errors are shown in Table IV. All the experiment videos can be found in the Supplementary Video.
@@
Based on the task 50, we also control the MSRA to bend 20 anticlockwise (50a20) and clockwise (50c20) on the x-y plane. The errors are shown in Table IV. We show the MSRA deformation on the same end position in Fig. 11(E), (F), and (G). The MSRA orientations on the x-y plane in these tasks are shown in Fig. 12. It is obvious that our approach can perform position and orientation control simultaneously and achieve satisfying accuracy even in some intricate tasks. The experimental results of the position and orientation control show that our strategy outperforms the PCC controller; hence, we employ it in the following high-level tasks.
IV-D Position Constraint
In these tasks, we aim to keep the end position of the base, middle, and end module invariant during motion, as depicted in Fig. 13. We include the losses $L_{p},L_{d},L_{o}$ for this task. In task ’base,’ we control the MSRA end to follow a circle trajectory while maintaining the base module position invariant, as shown in Fig. 13(A). In tasks ’middle’ and ’end,’ we aim to keep the middle and end module end positions invariant and change the same module end orientations from 0 to 360, as shown in Fig. 13(B) and (C). The errors are shown in Table V. During these motions, the corresponding module end positions remain static, but the orientations keep changing for MSRA motion. All the experiment videos can be found in the Supplementary Video.
TABLE V: Average errors and standard deviations in position constraint
@@ IV-D Position Constraint
IV-E Obstacle Avoidance
In obstacle avoidance, we include the losses $L_{p},L_{d}$ for position control and leverage the obstacle loss $L_{ob}$ ($u_{ob}=1$ in the following tasks). First, we set a red sponge bar as the target and control the MSRA end to reach it, as illustrated in Fig. 14(A). Then, we put one blue sponge bar as the obstacle in the last MSRA end trajectory and include $L_{ob}$ for replanning. The MSRA end first moves away from the obstacle to avoid collision (Fig. 14(B2)) and then reaches the target (Fig. 14(B3)). Furthermore, we set one more obstacle in the last MSRA end trajectory as depicted in Fig. 14(C). In this case, the S2C planning approach changes the trajectory and helps the MSRA reach the target.
In real-world applications such as domestic assistance, MSRA may encounter various obstacles, including humans, taps, fragile glasses, and lit gas stoves, which lead to different safety distances. By adjusting the obstacle threshold $r$ mentioned in Equation 5, the MSRA can show different performance on obstacles with different ’risk levels.’ As shown in Fig. 15, two obstacles are in the way of the MSRA end to the target. These obstacles are included in the planning optimization problem with two different obstacle losses $L_{ob}$, and the respective obstacle avoidance thresholds $r$ can represent their risk levels. If both obstacles are at the high-risk level, such as two fires on the stove, which means the obstacle threshold is large, the MSRA will vibrate near the obstacle boundaries instead of taking risks and rushing to the target, as shown in Fig. 15(B). However, if one obstacle is at low risk, like a tap, S2C will automatically generate available trajectories reaching the target, and the MSRA end will move to the target, as shown in Fig. 15(A) and (C). All these trajectories are planned by the optimization problem automatically, but they follow a basic rule: the MSRA end is expected to reach the target while being away from the high-risk obstacle. For instance, in Fig. 15(A), our strategy first leaves away from the right high-risk obstacle and then finds a path to reach the target between two obstacles. In Fig. 15(C), our strategy first leaves away from the left high-risk obstacle and then finds a path to reach the target on the right of the right obstacles. With these two kinds of trajectories, our strategy achieves the task of reaching the target while avoiding the obstacles according to their risk levels. All the experiment videos can be found in the Supplementary Video.
@@ IV-E Obstacle Avoidance
IV-F Online Interaction
If the obstacle and target will move during the task, it is necessary to plan the trajectory online. We include the losses $L_{p},L_{d},L_{ob}$ for position control and obstacle avoidance in this online task. During the motion, we cover the optical tracker markers with a red cloth to highlight that we achieve accurate control via only motor encoder feedback. As shown in Fig. 16(A), a blue obstacle is moved to reach the MSRA end by hand, and the MSRA end leaves the obstacle as long as it is inside the obstacle area, whose size depends on the obstacle threshold $r$. In addition to online obstacle avoidance, target following can also be fulfilled online. In Fig. 16(B), the MSRA end will follow the moving target. The MSRA end will stop if the target is out of the working space, but it will start to reach the target again if the target reenters the working space, as shown in Fig. 16(C). All the experiment videos can be found in the Supplementary Video.
The online target following motion is steady because the MSRA always aims to minimize the distance between the target and MSRA, which is motivated by the loss $L_{p}$. Meanwhile, in online obstacle avoidance, the MSRA motion is discontinuous. When the obstacle moves near the MSRA and their distance is smaller than $r$ in Equation 5, the MSRA will start to move away from the obstacle until the distance is larger than $r$. At this time, the MSRA will stop because of the configuration change loss $L_{d}$ preventing unnecessary motion, which leads to such discontinuous motion. MSRA will start to move when the obstacle catches the MSRA end again. The obstacle threshold $r$ constraints the obstacle avoidance $L_{ob}$ into a local area, hence the loss will only take effect when the MSRA is within the ’risk area,’ as the bordered blue circles in Fig. 15 and 16. Therefore, we endow the MSRA with situation awareness to some extent, which may be significant in the interaction tasks in future work.
@@ IV-F Online Interaction
V Conclusion and Discussion
This work proposes a versatile BiLSTM configuration space planning and control strategy S2C2A specifically for MSRAs flexible to various tasks. We train a biLSTM network $NN_{C2S}$ mapping from configuration to task space as the MSRA forward model and leverage it in an optimization problem S2C for configuration planning. Then, we train a biLSTM network $NN_{C2A}$ for configuration control C2A only using inaccurate internal sensing feedback. Utilizing a cable-driven MSRA, we demonstrate that our approach can outperform our previous PCC strategy via position and orientation control tasks. Furthermore, some challenging tasks have been managed, such as position constraints, obstacle avoidance, and online planning.
This flexible and versatile planning and control strategy endows MSRAs with accurate and complex deformation capability. Owing to modularity, MSRAs have the potential to perform sophisticated tasks, but effective control approaches specifically for MSRAs have not been fully explored. This work is a benchmark for MSRA control targeted at complicated deformation, enabling MSRA with flexibility and situation awareness. For instance, MSRAs can be utilized for minimally invasive surgery leveraging the position constraint ability and human interaction leveraging online planning ability.
There are some limitations to this strategy, which may be addressed in future work. First, it is challenging to estimate whether one target position is feasible for one MSRA, and the difficulty will increase when including target orientation additionally. Some physics-based approaches may be applied to MSRA working space construction. Also, this data-driven strategy may fail under an unknown payload or interaction. An online adjustment part like the kinematics controller in [10] is necessary to cope with this kind of online disturbance. Multiple internal sensors, such as IMU and flex sensors, may contribute to the accurate configuration estimation and feedback controller. Based on our data-driven strategy, we may either decrease the configuration estimation errors or apply the raw sensing feedback as a configuration without any physical models. Moreover, we would like to take full advantage of the situation awareness ability equipped by our strategy and perform some more intricate interaction tasks.